OpenStudy (anonymous):

Identity problem #2: sin^3 - 8/sin - 2 = sin^2 + 2sin + 4

OpenStudy (anonymous):

sin^4-8-2sin=sin^2+2sin+4.....

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

wait....nevermind these are proofs not equations...give me a minute.

OpenStudy (anonymous):

Okay, take your time. XD

OpenStudy (anonymous):

this is a tough one......

OpenStudy (anonymous):

they are fun though..give a minute....

OpenStudy (amistre64):

sin^3 - 8 -------- = sin^2 + 2sin + 4 ?? is this right? sin - 2

OpenStudy (amistre64):

the top is a difference of cubes.....

OpenStudy (amistre64):

(sin-2) (sin^2 +2sin +4) ------------------- (sin-2)

OpenStudy (anonymous):

really? i dont see that?

OpenStudy (amistre64):

sin^3 - 2^3.... plain as day :)

OpenStudy (anonymous):

yeah, but what about 1/sin?

OpenStudy (amistre64):

there is no 1/sin.... which problem you doin? :)

OpenStudy (anonymous):

same one, 8/sin

OpenStudy (amistre64):

nah.... that aint it.....

OpenStudy (amistre64):

(sin^3 - 8)/(sin - 2) = sin^2 + 2sin + 4

OpenStudy (anonymous):

there are no parentheses in the problem, so im assuming that it means -8/sin

OpenStudy (amistre64):

then it aint gonna work out :) Have a crack at it, but its futile.....

OpenStudy (anonymous):

It's (sin^3 - 8)/(sin-2)

OpenStudy (anonymous):

oh.....ok. You were right. My bad. sorry.

OpenStudy (amistre64):

'sok :) Blame it on the rain :)

OpenStudy (anonymous):

cool.

OpenStudy (anonymous):

How does a difference of cubes work? Can you show me it in detail?

OpenStudy (amistre64):

The best way to understand it is to actually multiply a cubed expression like (a-b)^3

OpenStudy (amistre64):

Which of course is just (a-b)^2 times (a-b)

OpenStudy (amistre64):

Ack!!...splained it wrong lol (a-b)(a^2 + ab + b^2)

OpenStudy (amistre64):

suppose you have an expression like this (f = first and l = last): (f^3 - l^3) can be factored: (f-l)(f^2 +fl +l^2)

OpenStudy (anonymous):

So what would be a and b? I'm sorry, I'm no good at this stuff. Thanks for helping though!

OpenStudy (amistre64):

Lets use some numbers: x^2 - 64 is a good example: 4*4*4 = 64 = 4^3 x^3 - 4^3 can be factored like this: (x-4)(x^2 +4x + 4^2)

OpenStudy (amistre64):

that should have been x^3 - 64....

OpenStudy (amistre64):

many centuries ago, mathmatickers noticed patterns in the way certain forms worked; and they surmised these tricks to help them solve the stuff quicker....

OpenStudy (anonymous):

Wow, I get it now! Thanks soooo much!!

OpenStudy (amistre64):

good job... I gotta be going home now.... Ciao :)

OpenStudy (anonymous):

Night!